# An Exploration of Knowledge Editing for Arabic

### v1

@@ Abstract
While Knowledge Editing (KE) has been widely explored in English, its behavior in morphologically rich languages like Arabic remains underexamined. In this work, we present the first study of Arabic KE. We evaluate four methods (ROME, MEMIT, ICE, and LTE) on Arabic translations of the ZsRE and Counterfact benchmarks, analyzing both multilingual and cross-lingual settings. Our experiments on Llama-2-7B-chat show show that parameter-based methods struggle with cross-lingual generalization, while instruction-tuned methods perform more robustly. We extend Learning-To-Edit (LTE) to a multilingual setting and show that joint Arabic-English training improves both editability and transfer. We release Arabic KE benchmarks and multilingual training for LTE data to support future research.
@@ 1 Introduction
Figure 1: Comparison of ROME, MEMIT, and ICE on LLaMA-7B-Chat across four metrics: reliability, generality, locality, and Portability
@@ 3.1 Data Curation
We use the NLLB-200 model111facebook/nllb-200-3.3B Team et al. (2022) to automatically translate ZsRE and Counterfact into Arabic and German. While synthetic, these translations are high-quality and provide the first large-scale KE benchmark for Arabic.
@@ 4.1 Arabic Editing Performance
Figure 1 compares four editing methods: ROME, MEMIT, ICE, and LTE-EN on Arabic edits using ZsRE dataset (Counterfact results are shown in 5 in Appendix A). LTE-EN consistently achieves the highest scores across reliability, generality, locality, and portability, indicating that instruction-tuned models, even when trained only on English, can generalize effectively to Arabic. ICE ranks second in reliability and generality, though its portability drops sharply on Counterfact, likely due to the challenge of counterfactual reasoning under zero-shot prompts. MEMIT excels in locality, preserving unrelated knowledge via its surgical update mechanism, but trails in generality and portability. ROME performs worst overall, highlighting the difficulty of transferring localized parameter edits to morphologically rich, non-English languages.
@@ 4.2 Multilingual Comparison
How does editing in Arabic compare with editing in other languages?
To assess cross-lingual robustness, we compare editing performance in Arabic, English, and German across four methods: ROME, MEMIT, ICE, and LTE-EN as shown in Figure 2 )(results for the counterfact dataset are shown in figure 6). Parameter-based methods (ROME and MEMIT) perform best in English but degrade noticeably in German and further in Arabic, reflecting their limited adaptability beyond English-tuned settings. In contrast, ICE exhibits stable performance across all three languages (Figure 2(c)), suggesting that prompt-based approaches are more resilient to linguistic variation. Similarly, LTE shows minimal degradation across languages, highlighting the benefits of instruction tuning for multilingual generalization.
@@
Figure 3: (a) Shows a comparison of the considered methods across the reliability,generality, locality, and portability metrics on the ZsRE dataset. (b) Shows a comparison of the averaged cross-lingual reliability scores on the ZsRE dataset and (c) Shows a comparison of the averaged cross-lingual generality scores on the ZsRE dataset. The x-axis in (b) and (c) refer to the language the edit is being applied in
@@ 5 Conclusion
We presented the first study of knowledge editing for Arabic, evaluating four editing paradigms: ROME, MEMIT, ICE, and LTE, on the ZsRE and Counterfact benchmarks. Our experiments reveal that parameter-based editing methods, though effective in English, struggle in Arabic and show poor crosslingual transfer. In contrast, instruction-tuned methods, especially our extended multilingual LTE framework, exhibit robust performance both in Arabic and across languages.
Our findings highlight key challenges and opportunities in multilingual knowledge editing. First, language-specific morphological and syntactic factors significantly affect the reliability and locality of edits. Second, crosslingual propagation is limited in most existing approaches, emphasizing the need for multilingual training. Finally, instruction tuning emerges as a promising direction for building language-agnostic editing capabilities.
We release benchmarks and multilingual training data to facilitate further research in this space. We hope this work serves as a foundation for future efforts aimed at scalable, reliable, and crosslingual knowledge editing for low-resource and morphologically rich languages like Arabic.
@@ Appendix A Additional Results
Figure 5: Comparison of ROME, MEMIT, and ICE on LLaMA-7B-Chat across four metrics: reliability, generality, locality, and Portability on the counterfact dataset

### v2

@@ Abstract
While Knowledge Editing (KE) has been widely explored in English, its behavior in morphologically rich languages like Arabic remains underexamined. In this work, we present the first study of Arabic KE. We evaluate four methods (ROME, MEMIT, ICE, and LTE) on Arabic translations of the ZsRE and Counterfact benchmarks, analyzing both multilingual and cross-lingual settings. Our experiments on Llama-2-7B-chat show that parameter-based methods struggle with cross-lingual generalization, while instruction-tuned methods perform more robustly. We extend Learning-To-Edit (LTE) to a multilingual setting and show that joint Arabic-English training improves both editability and transfer. We release Arabic KE benchmarks and multilingual training for LTE data to support future research.
@@ 1 Introduction
Figure 1: Comparison of ROME, MEMIT, and ICE on LLaMA2-7B-Chat across four metrics: reliability, generality, locality, and Portability
@@ 3.1 Data Curation
We use the NLLB-200 model [^1] https://hf.co/facebook/nllb-200-3.3B Team et al. (2022) to automatically translate ZsRE and Counterfact into Arabic and German. While synthetic, these translations are high-quality and provide the first large-scale KE benchmark for Arabic.  [^2] https://github.com/baselmousi/arabic-knowledge-editing
@@ 4.1 Arabic Editing Performance
Figure 1 compares four editing methods: ROME, MEMIT, ICE, and LTE-EN on Arabic edits using ZsRE dataset (Counterfact results are shown in figure 5 in Appendix A). LTE-EN consistently achieves the highest scores across reliability, generality, locality, and portability, indicating that instruction-tuned models, even when trained only on English, can generalize effectively to Arabic. ICE ranks second in reliability and generality, though its portability drops sharply on Counterfact, likely due to the challenge of counterfactual reasoning under zero-shot prompts. MEMIT excels in locality, preserving unrelated knowledge via its surgical update mechanism, but trails in generality and portability. ROME performs worst overall, highlighting the difficulty of transferring localized parameter edits to morphologically rich, non-English languages.
@@ 4.2 Multilingual Comparison
LLMs encode different languages in partially overlapping latent spaces Mousi et al. (2024). This raises an important research question: How does editing in Arabic compare to editing in other languages?
To assess cross-lingual robustness, we compare editing performance in Arabic, English, and German across four methods: ROME, MEMIT, ICE, and LTE-EN as shown in Figure 2 (Counterfact results are shown in figure 6). Parameter-based methods (ROME and MEMIT) perform best in English but degrade noticeably in German and further in Arabic, reflecting their limited adaptability beyond English-tuned settings. In contrast, ICE exhibits stable performance across all three languages (Figure 2(c)), suggesting that prompt-based approaches are more resilient to linguistic variation. Similarly, LTE shows minimal degradation across languages, highlighting the benefits of instruction tuning for multilingual generalization.
@@
Figure 3: (a) Shows a comparison of the considered methods across the reliability, generality, locality, and portability metrics on the ZsRE dataset. (b) Shows a comparison of the averaged cross-lingual reliability scores on the ZsRE dataset and (c) Shows a comparison of the averaged cross-lingual generality scores on the ZsRE dataset. The x-axis in (b) and (c) refer to the language the edit is being applied in.
@@ 5 Conclusion
We presented the first study of knowledge editing for Arabic, evaluating four editing paradigms: ROME, MEMIT, ICE, and LTE, on the ZsRE and Counterfact benchmarks. Our experiments reveal that parameter-based editing methods, though effective in English, struggle in Arabic and show poor crosslingual transfer. In contrast, instruction-tuned methods, especially our extended multilingual LTE framework, exhibit robust performance both in Arabic and across languages. Our findings highlight key challenges and opportunities in multilingual knowledge editing. First, language-specific morphological and syntactic factors significantly affect the reliability and locality of edits. Second, crosslingual propagation is limited in most existing approaches, emphasizing the need for multilingual training. Finally, instruction tuning emerges as a promising direction for building language-agnostic editing capabilities. We hope this work serves as a foundation for future efforts aimed at scalable and reliable knowledge editing for low-resource and morphologically rich languages like Arabic.
@@ Appendix A Additional Results
Figure 5: Comparison of ROME, MEMIT, and ICE on LLaMA2-7B-Chat across four metrics: reliability, generality, locality, and Portability on the counterfact dataset